OpenStudy (anonymous):

Help please!!! Solve the triangle,if possible.Round to the nearest hundredth B=52.8 degree C=114.6 degree b=10.13

OpenStudy (mathmale):

In this problem, two of the three angles of the triangle are given. What have you learned about the SUM of the interior angles of any triangle? Add angles B and C together. Then subtract this sum from 180 degrees (why?)

OpenStudy (anonymous):

to get A

OpenStudy (anonymous):

180-114.6-52.8=12.6 A=12.6

OpenStudy (mathmale):

Great! Now you have a triangle with three known angles and one known side. Which Law would be appropriate to use to determine each of the unknown sides, (a) the Law of Cosines, or (b) the Law of Sines?

OpenStudy (anonymous):

sin

OpenStudy (anonymous):

i'll use sin52.8/10.13 =a/sin12.6 right

OpenStudy (mathmale):

Right. Want to give it a try? sin A/a = sin B/b = sin C/c

OpenStudy (anonymous):

sin52.8/10.13=sin12.6/a

OpenStudy (anonymous):

what i have to do first

OpenStudy (mathmale):

Here you have an equation involving two ratios: sin 52.8 sin 12.6 ________ = _________ 10.13 a (It's easier to work with in this form.) I'd suggest you cross multiply. Then solve for a. Do the same to find side c, except use C = 114.6.

OpenStudy (mathmale):

If you use a calculator, be sure to set it to DEGREE mode first.

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

i got for a=2.77 c=11.56

OpenStudy (mathmale):

I haven't calculated those myself, but certainly would say that your results look very reasonable. Congrats! Any further questions?

OpenStudy (anonymous):

:) Thank you for your time

OpenStudy (mathmale):

My great pleasure. All the best to you.

OpenStudy (anonymous):

you too